Lauren plans to deposit $4000 into a bank account at the beginning of next month and $225/month into the same account at the end of that month and at the end of each subsequent month for the next 7 years. If her bank pays interest at a rate of 6%/year compounded monthly, how much will Lauren have in her account at the end of 7 years? (Assume she makes no withdrawals during the 7-year period. Round your answer to the nearest cent.)
Answer #1
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+6/(12*100))^12-1)*100
Effective Annual Rate% = 6.1678
Future value = present value*(1+ rate)^time
Future value = 4000*(1+0.061678)^7
Future value = 6081.49
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 225*(((1+ 6.1678/1200)^(7*12)-1)/(6.1678/1200))
FV = 23561.94

Total = 6081.49+23561.94

=

29643.43